if | A| = -7 find x where A ( 3 X 2 1)

1
Expert's answer
2021-12-08T12:44:38-0500
"A=\begin{pmatrix}
 3 & x \\
 2 & 1
\end{pmatrix}"

"|A|=\begin{vmatrix}
 3 & x \\
 2 & 1
\end{vmatrix}=3(1)-2x=3-2x=-7"

"2x=10"

"x=5"